### v1

# To cover a permutohedron

Abstract
The permutohedron $P_{n}$ of order $n$ is a polytope embedded in $\mathbb{R}^{n}$ whose vertex coordinates are permutations of the first $n$ natural numbers. It is obvious that $P_{n}$ lies on the hyperplane $H_{n}$ consisting of points whose coordinates sum up to $n(n+1)/2$. We prove that if the vertices of $P_{n}$ are contained in the union of $m$ affine hyperplanes different from $H_{n}$, then $m\geq n$ when $n\geq 3$ is odd, and $m\geq n-1$ when $n\geq 4$ is even. This confirms a recent conjecture of Hegedüs and Károlyi. Our proof gives an algebraic criterion for a non-standard permutohedron generated by $n$ distinct real numbers to require at least $n$ non-trivial hyperplanes to cover its vertices.
Let $A$ be a set of $n$ distinct real numbers. We use $P_{A}$ to denote the polytope in $\mathbb{R}^{n}$ whose vertex coordinates are permutations of $A$. It is easy to argue that all these points, whose coordinates are permutations of $A$, are in convex position. It is also obvious that $P_{A}$ is contained in the hyperplane $H_{A}$ defined by the equation $\sum_{j\in[n]}x_{j}=\sum_{a\in A}a$. Here, $n]:=\{1,2,\dots,n\}$ and $x_{j}$ is the $j$-th coordinate. In fact, $P_{A}$ is always (n1)(n-1)-dimensional, see, e.g. [5]. The special case $P_{[n]}$ is known as the permutohedron of order $n$. We write $P_{n}:=P_{[n]}$ and $H_{n}:=H_{[n]}$ for simplicity.
A collection $\mathcal{C}$ of affine hyperplanes is called a vertex cover of $P_{A}$ if $H_{A}\not\in\mathcal{C}$ and every vertex of $P_{A}$ lies on some hyperplane in $\mathcal{C}$. It is obvious that $P_{A}$ can always be covered by $n$ hyperplanes defined by equations $x_{1}=a$ for $a\in A$. However, when $n$ is even, there are n1n-1 hyperplanes, defined by $x_{1}+x_{j}=n+1$ for $j\in[n]\setminus 1$, that contain all vertices of $P_{n}$. Recently, Hegedüs and Károlyi [4] conjectured the following statement, which is our main result.
Theorem 1.
If $n\geq 3$ is odd, a vertex cover of $P_{n}$ must have size at least $n$. If $n\geq 4$ is even, a vertex cover of $P_{n}$ must have size at least n1n-1.
Our proof gives an algebraic criterion on $A$ ensuring that any vertex cover of $P_{A}$ has size at least $n$. We write the elementary symmetric polynomial on $n$ variables of degree $d$ as
$S_{d}(\textbf{x})=S_{d}(x_{1},\dots,x_{n})=\sum_{1\leq j_{1}<j_{2}<\dots<j_{d}\leq n}x_{j_{1}}x_{j_{2}}\cdots x_{j_{d}}.$
By abuse of notation, we let $S_{d}(A)$ be the value of $S_{d}$ at a point whose coordinate is any permutation of $A$. We consider the following polynomial in one complex variable
$F_{A}(z)=z^{n}-S_{1}(A)\cdot z^{n-1}+S_{2}(A)\cdot z^{n-2}-\cdots+(-1)^{n-1}S_{n-1}(A)\cdot z.$
A critical point of $F_{A}$ refers to a number $p$ such that $F^{\prime}_{A}(p)=0$. We define a critical value of $F_{A}$ to be a number $v$ such that $F_{A}(p)+(-1)^{n}v=0$ for some critical point $p$. By elementary algebra, $v$ is a critical value if and only if the equation $F_{A}(z)+(-1)^{n}v=0$ has a multiple root. By elementary calculus, $F_{A}$ actually has n1n-1 distinct real critical points interlaced between the consecutive elements in $A$. Hence, all critical values of $F_{A}$ are real numbers, though not necessarily distinct. The importance of complex numbers will be evident in the proof of our criterion.
Theorem 2.
If there are n1n-1 distinct critical values of $F_{A}$, then a vertex cover of $P_{A}$ must have size at least $n$.
We first deduce Theorem 1 from Theorem 2. Moreover, as a consequence of Theorem 2, $P_{A}$ generated by a generic $A$ would require at least $n$ elements in its vertex cover.
Proof of Theorem 1.
Let $n\geq 3$ be odd and write $F_{n}=F_{[n]}$. According to Theorem 2, it suffices to verify that $F_{n}$ has n1n-1 distinct critical values. We consider the real polynomial $G_{n}(x)=\prod_{j\in[n]}(x-j)$ and notice that $G_{n}(x)=F_{n}(x)-n!$. Hence, the critical points of $G_{n}$ coincide with those of $F_{n}$, and it suffices to prove $G_{n}$ has n1n-1 distinct critical values.
Now, let $v_{j}$ be the extreme value of $G_{n}$ on the interval $(j,j+1)$ for $j\in[n-1]$. Notice that $v_{1},\dots,v_{n-1}$ are the critical values of $G_{n}$. Since $n$ is odd, the graph of the function $G_{n}$ is symmetric with respect to the point $\frac{n+1}{2}$ on the $x$-axis. Therefore, it suffices to prove $|v_{j}|<|v_{j+1}|$ for every integer $\frac{n}{2}<j<n$. Let $\delta$ be defined by $v_{j}=G_{n}(j+1-\delta)$, we can compute
$\frac{|G_{n}(j+1-\delta)|}{|G_{n}(j+1+\delta)|}$ $=\frac{(j-\delta)(j-1-\delta)\cdots(1-\delta)\cdot\delta(1+\delta)\cdots(n-j-1+\delta)}{(j+\delta)(j-1+\delta)\cdots\delta\cdot(1-\delta)(2-\delta)\cdots(n-j-1-\delta)}$
$=\frac{(n-j-\delta)(n-j-1-\delta)\cdots(j-\delta)}{(n-j+\delta)(n-j-1+\delta)\cdots(j+\delta)}<1.$
This implies $|v_{j}|=|G_{n}(j+1-\delta)|<|G_{n}(j+1+\delta)|\leq|v_{j+1}|$ as wanted.
Let $n\geq 4$ be even. We use $Q_{j}$ to denote the intersection between $P_{n}$ and the hyperplane defined by $x_{j}=n$ for $j\in[n]$. Notice that the affine span of $Q_{j}\cup Q_{k}$ ($j\neq k$) has dimension n1n-1, which contains the whole $P_{n}$. Hence an affine hyperplane different from $H_{n}$ can fully contain at most one $Q_{j}$. Suppose $P_{n}$ has a vertex cover of size less than n1n-1. Then there exists $k$ such that $Q_{k}$ is not contained in any hyperplane from this cover. On the other hand, $Q_{k}$ can be regarded as a copy of $P_{n-1}$, so we have found a vertex cover of $P_{n-1}$ of size less than n1n-1, contradicting our conclusion for the case when $n\geq 3$ is odd. ∎
Our proof of Theorem 2 requires surface-level knowledge of complex algebraic geometry. We refer the reader to the first section of [2] and the first chapter of [1] for an introduction. Given a covering map $f:X\to Y$ and a base point $y\in Y$, the monodromy action of the fundamental group $\pi_{1}(Y,y)$ on the fiber $f^{-1}(y)$ is as follows: we can always lift a loop representing $\ell\in\pi_{1}(Y,y)$ starting at a point $x\in f^{-1}(y)$; the image of $x$ under the action of $\ell$ is the ending point of this lifting, and the monodromy theorem guarantees this ending point is well-defined. To avoid wordiness in our proof, “near” a point means “on some neighborhood of” that point.
Proof of Theorem 2.
Let $V\subset\mathbb{C}^{n}$ be the analytic variety (not necessarily irreducible) defined by the vanishing of polynomials $S_{d}(\textbf{x})-S_{d}(A)$ for $d\in[n-1]$. Note that all vertices of $P_{A}$ are on $V$. It suffices to prove that $V$ is irreducible and one-dimensional (over $\mathbb{C}$). Indeed, a hyperplane $H$ in $\mathbb{R}^{n}$ can also be regarded as a hyperplane in $\mathbb{C}^{n}$. Given that $V$ is an irreducible curve, $V\cap H$ either equals $V$ or has dimension zero. In the former case, all vertices of $P_{A}$ are contained in $H$, which implies $H=H_{A}$. In the latter case, we have $|V\cap H|\leq(n-1)!$ by Bézout’s theorem. Hence, a hyperplane different from $H_{A}$ contains at most (n1)!(n-1)! vertices of $P_{A}$. Since $P_{A}$ has $n!$ vertices, its vertex cover must have size at least $n$.
We consider the holomorphic function $f:V\to\mathbb{C}$ such that $f(x_{1},x_{2},\dots,x_{n})=x_{1}x_{2}\dots x_{n}$. Let $R_{y}$ be the collection of $n$ roots (possibly repeated) of the equation $F_{A}(z)+(-1)^{n}y=0$ for fixed $y\in\mathbb{C}$. Importantly, the preimage $f^{-1}(y)$ consists of all permutations of $R_{y}$. In particular, $|f^{-1}(y)|<\infty$ for all $y\in\mathbb{C}$. This means any irreducible component of $V$ has dimension at most one. Because the roots of a polynomial vary continuously as a function of the coefficients (see e.g. [3]), $V$ does not have isolated points, so $V$ is purely one-dimensional. It suffices to prove $V$ is irreducible.
Now, we write $\Omega=\mathbb{C}\setminus\{v_{1},v_{2},\dots,v_{n-1}\}$ with $v_{1},\dots,v_{n-1}$ being the critical values of $F_{A}$. Let $U=f^{-1}(\Omega)$. We can regard $f:U\to\Omega$ as a covering map between Riemann surfaces (not necessarily connected) as follows: for a point $\textbf{x}\in U$, there are holomorphic functions $x_{1},\dots,x_{n}$ defined near $f(\textbf{x})$ with $(x_{1}(f(\textbf{x})),\dots,x_{n}(f(\textbf{x})))=\textbf{x}$; moreover, $F_{A}(x_{j}(y))+(-1)^{n}y=0$ for all $y$ near $f(\textbf{x})$, see e.g. Corollary 8.8 in [1]; then the holomorphic mapping $(x_{1}(y),\dots,x_{n}(y))$ near $f(\textbf{x})$ is a parametrization of $U$ near x. Since $V$ is purely one-dimensional and $U$ differs from $V$ by finitely many points, $U$ being irreducible implies $V$ being irreducible. Since $U$ is smooth by its parametrization, its irreducibility is implied by its connectedness. As $\Omega$ is path-connected, we only need to show points in $f^{-1}(y)$ are connected by paths in $U$ for some $y\in\Omega$.
It suffices to prove the monodromy action of $\pi_{1}(\Omega,y)$ on the fiber $f^{-1}(y)$ is transitive for some $y\in\Omega$. Note that a permutation of $R_{y}$ naturally acts on $f^{-1}(y)$ by permuting the coordinates. We have the following two claims.
Claim 3.
For $y\in\Omega$ and $j\in[n-1]$, let $\ell_{j}\in\pi_{1}(\Omega,y)$ be represented by a loop at $y$ whose winding number equals 1 around $v_{j}$, and equals 0 around $v_{k}$ for $k\neq j$. Then the monodromy action of $\ell_{j}$ on $f^{-1}(y)$ is the same as a transposition in the permutation group of $R_{y}$.
Claim 4.
For a complex number $y$ with sufficiently large $|y|$, let $\ell_{n}\in\pi_{1}(\Omega,y)$ be represented by a loop at $y$ whose trajectory is a circle centered at origin. Then the monodromy action of $\ell_{n}$ on $f^{-1}(y)$ is the same as a cycle of length $n$ in the permutation group of $R_{y}$.
We choose the loops such that $\ell_{1}\cdot\ell_{2}\cdots\ell_{n-1}=\ell_{n}$ in $\pi_{1}(\Omega,y)$ as in Figure 1. On the other hand, in a permutation group of $n$ elements, if an $n$-cycle factors into n1n-1 transpositions, then they will generate the whole group. Therefore, we have all permutations of $R_{y}$ acting on $f^{-1}(y)$ in the monodromy action, which is the transitivity as we wanted.
Figure 1: $\ell_{1}\cdot\ell_{2}\cdots\ell_{n-1}=\ell_{n}$ in the fundamental group.
Next, we give a proof for Claim 3. By our hypothesis, $F_{A}$ has n1n-1 distinct critical points $p_{1},\dots,p_{n-1}$. Crucially, every $p_{j}$ must be a simple critical point, that is, $F_{A}^{\prime\prime}(p_{j})\neq 0$. We consider the multi-variable complex function $\alpha(z,y):=F_{A}(z)+(-1)^{n}y$. We can compute
$\frac{\partial\alpha}{\partial z}(p_{j},v_{j})=F^{\prime}_{A}(p_{j})=0\quad\text{and}\quad\frac{\partial^{2}\alpha}{\partial z^{2}}(p_{j},v_{j})=F^{\prime\prime}_{A}(p_{j})\neq 0.$
According to the Weierstrass preparation theorem, near $(p_{j},v_{j})$, we can write
$\alpha(z,y)=\left((z-p_{j})^{2}-\beta(y)\right)\cdot\gamma(z,y),$
for holomorphic functions $\beta,\gamma$ such that $\beta(v_{j})=0$ and $\gamma(z,y)\neq 0$. We also have $\beta^{\prime}(v_{j})\neq 0$ by computing $\partial\alpha/\partial y$. This implies $\beta$ is non-zero near $v_{j}$ except at $v_{j}$, and cannot be the square of a holomorphic function near $v_{j}$. For a point $y_{\ast}$ near but not equal to $v_{j}$, there is a holomorphic function $\delta$ defined near $y_{\ast}$ with $\delta^{2}=\beta$ (see e.g. Lemma 8.7 in [1]), then $z_{\pm}(y)=p_{j}\pm\delta(y)$ are two functions satisfying $F_{A}(z_{\pm}(y))+(-1)^{n}y=0$ near $y_{\ast}$. Hence, $z_{+}$ and $z_{-}$ are two coordinate functions in the parametrization of $U$ near $y_{\ast}$. It is a standard argument, see Example 8.10 in [1], that the analytic continuation of a small loop around $v_{j}$ transposes $z_{+}$ and $z_{-}$.
Finally, we give a proof for Claim 4. To this end, we consider change of coordinates $s=1/z$ and $h=1/y$. Let $\epsilon(s)=(-1)^{n-1}(1/s^{n})(1/F_{A}(1/s))$. We can remove the singularity of $\epsilon$ at $s=0$ by defining $\epsilon(0)=(-1)^{n-1}$. There is a holomorphic function $\zeta$ defined near $s=0$ with $\zeta^{n}=\epsilon$. We write $\eta(s)=s\cdot\zeta(s)$ and compute $\eta^{\prime}(0)\neq 0$, which means $\eta$ is invertible at $s=0$. Observe that $|y|\to\infty$ implies $|z|\to\infty$ under the condition $F_{A}(z)+(-1)^{n}y=0$. Hence, near $h=0$, we have
$F_{A}(z)+(-1)^{n}y=0\iff s^{n}\cdot\epsilon(s)=h\iff(\eta(s))^{n}=h.$
For a point $h_{\ast}$ near but not equal to $h=0$, there is a holomorphic function $\theta$ defined near $h_{\ast}$ such that $\theta^{n}=h$, then $s_{j}(h)=\eta^{-1}(\iota^{j}\theta(h))$ is a function satisfying $(\eta(s_{j}(h)))^{n}=h$ for $j\in[n]$. Here, $\iota=\exp(2\pi i/n)$ is the $n$-th root of unity. Hence, $1/s_{1},\dots,1/s_{n}$ are the coordinate functions in the parametrization of $U$ near $1/h_{\ast}$. It is a standard argument that the analytic continuation along a small loop around $h=0$ takes $s_{j}$ to $s_{j+1}$, and a small loop around $h=0$ is a large loop around $y=0$, which is homotopic to $\ell_{n}$. ∎
Let us baldly propose the conjecture below. Following our argument for the even case of Theorem 1, this conjecture implies that a vertex cover of $P_{A}$ must have size at least |A|1|A|-1 for any set $A$ of distinct real numbers.
Conjecture 5.
If the size of a set $A$ of distinct real numbers is odd, then a vertex cover of $P_{A}$ must have size at least $|A|$.
Acknowledgement. Theorem 1 was communicated to Ji Zeng as a conjecture by Gyula Károlyi at the 18-th Emléktábla workshop in 2025. We wish to thank Nóra Frankl, Gyula Károlyi, Gergely Kiss, and Gábor Somlai for efforts and discussions on this problem. We also wish to thank Zoltán Lóránt Nagy for organizing the workshop.